Lock, Statistics: Unlocking the Power of Data, 2e Help I S information for all players who averaged more than 24 minutes per game, and includes n = 82 players and 4 variables. Which Accomplishment of LeBron James is Most Impressive? i -trt bel и st we th. ne an ar stand de al ins or our the va ables in t e NBA players2015 dataset. FGPt is the field goal percentage ponts is total number of points 0.488, scored 1743 points, had 511 assists, and had 109 steals. The dataset NBAPlayers2015 contains information on many variables for players in the NBA (National Basketball Association) during the 2014-2015 season. The dataset indudes s total number of assists during the season, and Steals is total number of steals during the season. LeBron James had a field goal percentage of Variable 0.453 915 205 67.5 0.054 357 149 33.6 Steals Table 1 Summary statistics on NBA players Find the z-score for each of LeBrons statistics Round your answers to three decimal places score for 5:18 PM e here to search
Answer #1

Here we have :

Z score= (x-\mu)/\sigma

z score FG = (0.488 - 0.453)/0.054 = 0.648

z score points= (1743 - 915)/357 = 2.319

z score assists = (511 - 205)/149= 2.054

z score steals = (109 - 67.5)/33.6 = 1.235